News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 19 January 2001 8.00am EST

Last week the observing schedule was re-planned to include a Target Of Opportunity observation of the dwarf nova SS Cygni that is currently entering an outburst state. The observation was completed successfully following the insertion of the LETG during a real-time pass on Jan 17. Planned observations of IC1727, A262 and NGC266 were impacted by the TOO and will be rescheduled.

Other real-time procedures last week included a check of the CPE warm start count that showed the value remained at 1.

Of note last week was the conduct of a training simulation for the full operation team. The simulation was run over 2 days using the spacecraft hardware simulator as the source of data. The team had no prior knowledge of the simulated problem (a failed gyro leading to safemode) and responded with a correct analysis, problem identification and plan for recovery to science operations. The exercise also included several simulated ground system problems.

The planned observations for the coming week are shown below and include the retraction of the LETG on Jan 24.

------------------------------------------
H1821+643               ACIS-S/LETG  Jan 23 
HER X-1                 HRC-S/LETG   Jan 24 
RX J0822-4300           HRC-S        Jan 25 
2MASSJ034857+1255       ACIS-I
2MASSJ023430+2438       ACIS-I
2MASSJ005055+2933       ACIS-I
AGX J233200+1945        ACIS-I
Radiation Belts
2MASSJ010835+2148       ACIS-I       Jan 26 
2MASSJ012031+2003       ACIS-I
2MASSJ015721+1712       ACIS-I
CADIS 01H Field         ACIS-I
A 478                   ACIS-S
3C120                   ACIS-S       Jan 27 
NGC 7331                ACIS-S
NGC 660                 ACIS-S
NGC 741 Group           ACIS-S
1RXSJ0203-2438          HRC-S        Jan 28 
NGC1097                 ACIS-S
Radiation Belts
MACS J0717.5+3745       ACIS-I       Jan 29 
2MASSJ095504+1705       ACIS-I
MS0839.9+2938           ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports